after being discounted 30%, a weather radio sells for $62.97 find the original price?

Answer by stanbon(75887)

after being discounted 30%, a weather radio sells for $62.97 find the original price?
-----
x - 0.30x = 62.97
0.7x = 62.97
x = 62.97/0.7 = $89.96
----------
Cheers,
Stan H.
----------